# Memory Self-Regeneration: Uncovering Hidden Knowledge in Unlearned Models

### v1

@@ Abstract
The impressive capability of modern text-to-image models to generate realistic visuals has come with a serious drawback: they can be misused to create harmful, deceptive or unlawful content. This has accelerated the push for machine unlearning. This new field seeks to selectively remove specific knowledge from a model’s training data without causing a drop in its overall performance. However, it turns out that actually forgetting a given concept is an extremely difficult task. Models exposed to attacks using adversarial prompts show the ability to generate so-called unlearned concepts, which can be not only harmful but also illegal. In this paper, we present considerations regarding the ability of models to forget and recall knowledge, introducing the Memory Self-Regeneration task. Furthermore, we present MemoRa strategy, which we consider to be a regenerative approach supporting the effective recovery of previously lost knowledge. Moreover, we propose that robustness in knowledge retrieval is a crucial yet underexplored evaluation measure for developing more robust and effective unlearning techniques. Finally, we demonstrate that forgetting occurs in two distinct ways: short-term, where concepts can be quickly recalled, and long-term, where recovery is more challenging. Code is available at https://github.com/gmum/MemoRa.
@@ 1 Introduction
Figure 2: Our method aims to recover unlearned information by using only a few images containing removed concepts. We first expand the training set using DDIM inversion and diversify it via spherical interpolation. Next, we fine-tune a LoRA adapter to restore the erased concept. Results reveal two types of forgetting: short-term, where knowledge is quickly recovered, and long-term, where recovery is harder. We hypothesize that short-term forgetting corresponds to parts of the manifold moving away (the class is erased, but other classes show lower FID), while long-term forgetting reflects a displacement along the manifold.
@@
To investigate this phenomenon, we propose the MemoRa (Memory Regeneration with LoRA) strategy, which we consider a regenerative process, showing how knowledge recovery affects unlearning models. An overview of the strategy is presented in Fig. 2. Our method begins by applying DDIM inversion to reconstruct latent trajectories of images corresponding to the removed concept. To overcome the scarcity of available samples, we expand this dataset using spherical interpolation in latent space, which provides diverse yet consistent training examples. Next, instead of fine-tuning the full model, we update only a lightweight LoRA adapter, enabling efficient retraining with minimal computational cost. This design makes MemoRa practical even under resource constraints, while also serving as a diagnostic tool to probe the depth of forgetting. We demonstrate that some models fail to truly forget the targeted concept. Furthermore, we demonstrate that certain models exhibit a particular predisposition to rapidly recall previously forgotten concepts, indicating that this information is still stored in structures analogous to human long-term memory. In contrast, models with lower regenerative capacity show greater memory loss and more effective unlearning.
Specific models are able to rapidly recover the erased knowledge, while others require extensive fine-tuning. We hypothesize that this difference arises from the way unlearning methods affect the underlying representation manifold. In particular, certain approaches result in a sudden drop in the FID score across concepts, indicating that the model has moved outside the manifold, leading to shallow, short-term forgetting that can be quickly reversed. In contrast, methods associated with long-term forgetting tend to preserve FID scores across other classes, which indicates that the model remains within the manifold but replaces the erased knowledge with alternative representations, consequently making recovery more difficult.
@@ 2 Related Works
More recent work on unlearning within diffusion models focuses on parameter updates that suppress unwanted concepts. EDiff (Wu et al., 2024) achieves this through a bi-level optimization, while ESD (Gandikota et al., 2023) modifies classifier-free guidance by incorporating negative prompts. FMN (Zhang et al., 2024a) introduces a targeted loss on attention layers to steer the forgetting process. SalUn (Fan et al., 2023) and SHS (Wu & Harandi, 2024) adapt model weights by exploiting saliency and sensitivity analyses to localize parameters relevant to the concept. SEMU (Sendera et al., 2025) leverages Singular Value Decomposition (SVD) to project representations into a lower-dimensional space that facilitates selective erasure. Selective Ablation (SA) (Heng & Soh, 2023) replaces the distribution of a forbidden concept with that of a surrogate, an idea extended in Concept Ablation (CA) (Kumari et al., 2023) using predefined anchors. In another direction, SPM (Lyu et al., 2024) employs structural interventions by inserting lightweight linear adapters that block the flow of targeted features. SAeUron (Cywiński & Deja, 2025) applies sparse autoencoders to isolate and suppress concept-specific representations, offering interpretable and robust unlearning with minimal performance degradation, even under adversarial prompting. On the other hand, AdvUnlearn (Zhang et al., 2024c) proposes the use of adversarial training to directly optimize the text encoder.
@@
There are various techniques for attacking DMs to bypass protections against malicious content generation. The basic methods involve manipulating words, which includes replacing, deleting letters, inserting additional characters, and paraphrasing an unlearned concept (Eger & Benz, 2020; Li et al., 2018; Garg & Ramakrishnan, 2020). Other attack techniques include textual inversion (Gal et al., 2022), utilizing a frozen U-Net as a guide (white box) (Chin et al., 2023), or UnlearnDiffAtk (Zhang et al., 2024b), which generates adversarial prompts without needing an auxiliary model. The approaches discussed above enable the removal of targeted concepts from text-to-image models. Yet, regardless of the strategy employed, an important question remains: to what extent is the forgetting actually effective? Evidence suggests that current methods have notable limitations, as erased concepts can often be recovered. For instance, UnlearnDiffAtk (Zhang et al., 2024b) shows that both prompt tuning and backdooring techniques (Wang et al., 2024; Grebe et al., 2025) can restore removed concepts. Example images obtained from adversarial attacks are shown in Fig. 3. A key aspect of our pipeline, consistent with the methods discussed above, is the assumption of access to the model. It is worth noting that prompt tuning is computationally demanding, requiring separate attacks for each prompt, and fine-tuning alters concept representations in ways that make it difficult to pinpoint which concepts are affected. These studies highlight the inherent shortcomings of existing unlearning methods.
In this paper, we demonstrate that this problem is more complex and has significant implications. We highlight that concepts can be revisited and relearned using only a limited number of images.
@@ 3 How to Relearn diffusion models using MemoRa strategy
Standard evaluation metrics primarily utilize a scheme for generating images from prompts that use words related to the unlearned phrase, paraphrase the concept, or use synonyms. This approach only analyzes the model’s performance in user-tested mode, but does not examine which representations remain and which are removed from the latent space. Fig. 4 displays the comprehensive inversion of the entire photo denoising process for two unlearned techniques. It investigates how different methods respond to encoding a prohibited image and evaluates the effectiveness of these methods in erasing the semantic features associated with the forbidden concept in the latent representation after inversion. If the images of this concept reproduce, we are dealing with short-term forgetting, while if the inversion process leads to deviated latencies relative to the original versions, we are dealing with long-term memory forgetting. It is noticeable that the ESD method (Gandikota et al., 2023) is sensitive to repeated reconstruction of the deleted concept, whereas MACE (Lu et al., 2024) is a more aggressive approach that removes this representation more efficiently.
@@
Additionally, we use the extracted latents to generate more images by spherical interpolation. Spherical interpolation is a complement to Stage 1, as illustrated in Fig. 2, and determines intermediate values between two established latents:
$\text{slerp}\!\left(z_{T}^{(1)},z_{T}^{(2)};p\right)=\frac{\sin\!\big((1-p)\Omega\big)}{\sin\Omega}\,z_{T}^{(1)}+\frac{\sin\!\big(p\Omega\big)}{\sin\Omega}\,z_{T}^{(2)},\qquad p\in[0,1].$
where $p$ is the hyperparameter of interpolation and the $\Omega$ is calculated by:
$\Omega=\arccos\!\left(\frac{\langle z_{T}^{(1)},z_{T}^{(2)}\rangle}{\|z_{T}^{(1)}\|\,\|z_{T}^{(2)}\|}\right).$
@@
4 Experiments
In this section, we present an evaluation of MemoRa for restoring knowledge across three main categories: nudity, objects, and styles. For this purpose, we utilized publicly available weights from unlearned models from Zhang et al. (2024c), which includes SOTA unlearning models. MemoRa is a strategy that stands out for its simplicity, making it easily applicable to all the aforementioned methods. Remarkably, MemoRa enables these models to recover lost knowledge in a much faster and straightforward manner. Additional experiments are presented in the Appendix.
Dataset Metrics SD v1.4 FMN SPM ESD
base unlearn MemoRa unlearn MemoRa unlearn MemoRa
CLIP $(\uparrow)$ 31.08 30.69 31.21 31.07 31.34 29.95 30.84
Dataset Metrics AdvUnlearn SalUn ED SH
unlearn MemoRa unlearn MemoRa unlearn MemoRa unlearn MemoRa
CLIP $(\uparrow)$ 30.56 30.44 31.13 31.16 30.88 31.00 27.72 27.68
@@ 4 Experiments
Evaluation Setups To assess the effectiveness and performance of MemoRa, we conduct a similar evaluation as in (Zhang et al., 2024c), which also includes testing the models’ robustness to attacks using adversarial prompts, employing the UnlearnDiffAtk technique.
@@
To evaluate MemoRa, we use a complex measure: the attack success rate (ASR) (Zhang et al., 2024b). To calculate effectiveness, a set of 50 prompts generated with GPT-4 was used, which contained target training words from the Imagenette dataset. These prompts were previously tested to ensure that the original Stable Diffusion 1.4 could generate correct images from them. The ASR metric can be divided into two components: the pre-attack success rate (pre-ASR) and the post-attack success rate (post-ASR). The pre-ASR metric reflects the model’s unlearned knowledge, as it is evaluated under normal conditions without any attacks. In contrast, the post-ASR metric assesses the effectiveness of attacks on the unlearned model using adversarial prompts. These metrics are not correlated, as high unlearned data will not always lead to low attack effectiveness. Therefore, for the purposes of our experiment, we consider these two metrics separately. The "No Attack" scenario will represent the pre-ASR, and "UnlearnDiffAtk" will encompass both pre-ASR and post-ASR to allow for a more thorough assessment of the relearning of the models mentioned. The higher the "No Attack" measure, the better the model generates images in default mode. Conversely, a higher "UnlearnDiffAtk" measure indicates lower resistance to attacks. We used ImageNet-pretrained ResNet-50 to classify images for attack evaluation.
@@
Nudity Relearning
@@
Table 1 presents the results using the MemoRa strategy. Each method experienced an increase in knowledge, but at varying levels. MemoRa suggests that SPM and ESD are methods in which unlearning is shallow and a large return to pre-unlearning knowledge is possible. MemoRa allowed for the unlearned knowledge to be recovered, as symbolized by the significantly increasing No Attack measure. Fig. 5 compares all unlearning methods that used the MemoRa strategy. The SH, SalUn, and ED methods remove nudity features the most in the generated images even after applying the memory regeneration strategy. The disadvantage of these approaches is the deterioration of the remaining knowledge, where MemoRa corrects it closer to the original state. Sample visualizations of this situation are presented in Fig. 6.
Fig. 7 presents a comparison between concept recovery for UnlearnDiffAtk and MemoRa. Both strategies add the forbidden concept, but the images for MemoRa contain more nudity features in individual photos. In particular the Belly, Female Genitals and Male Breast. UnlearnDiffatk aims to generate at least one forbidden feature, while LoRA applies the nudity concept overall, creating photos more similar to the original ones. More visual and quantitative results for MemoRa are provided in the Appendix B.1.
Objects Relearning Three classes from the Imagenette (Shleifer & Prokop, 2019) dataset were used: parachute, church and garbage truck. Tab. 2 referring to unlearning-relearning concept parachute. We can observe the weakest unlearning for FMN, SPM. Applying MemoRa sheds light on the actual forgetting, revealing that some methods achieve high recall rates in an instant, these are: FMN, SPM, ESD, SalUn, and ED. Forgotten knowledge about concept can be recovered to approximately 85% from just a few percent and even 100% after using additional attacks. In summary, more effective unlearning does not always translate into greater resistance to knowledge retrieval.
@@
In Fig. 8 the results of using MemoRa to restore the church are presented, where the effectiveness of the strategy is noticeable. MemoRa can recreate the original photo even for methods that significantly changed the trajectory for the unlearned concept. MemoRa increases the number of successful attacks, leading to shorter inference times using adversarial prompts, see Tab. 3. Introducing attacking involves time-consuming inference, whereas with MemoRa + UnlearnDiffAtk strategy, inference time is much shorter (for example from 10 minutes to 4 minutes). MemoRa requires only pre-training (approximately 15 minutes), and its image generation time matches that of the base model.. More visual and quantitative results are provided in the Appendix B.2.
Style Relearning For style evaluation, prompts were used according to the setup described in (Gandikota et al., 2023), and the ViT-base model, fine-tuned on the WikiArt dataset, was used as the classifier. However, the classifier may struggle to correctly classify such paintings, as the paintings are very intense and slightly deviate from the original. Nevertheless, we consider them examples of Van Gogh’s style because the main characteristic features are clearly present.
Fig. 9 presents a comparison between images before and after the MemoRa strategy. Images generated after applying the reminder strategy demonstrate the distinctive style of Vincent Van Gogh. MemoRa accurately replicated the Starry Night scene, incorporating strong brushstrokes, swirls in the sky, and a contrast of blues and yellows. Therefore, it seems that the models do not long-term forget about this distinctive style. More numerical and visual results are provided in Appendix B.3.
Trade-off Between Re-Learning and Utility During inference, the LoRA adapter for all tasks causes the FID metric to increase compared to the unlearned model, while the CLIP metric remains high, as shown in Tab. 1 and Tab. 2. To address this issue, we employ the Autoguidance technique (Kasymov et al., 2024; Karras et al., 2024), which involves using a combination (interpolation) of responses from both models during generation, rather than relying on just one model. The setup for AutoMemoRa is described in Appendix A.
Multi-MemoRa Often, a model that has undergone unlearning may forget multiple concepts. We demonstrate the use of multi-MemoRa on the task of relearning well-known celebrities. Combining adapters and recalling multiple items for our strategy is not difficult. The images in Fig. 11 demonstrate that the MACE model only shallowly unlearned the selected celebrities, as Multi-MemoRa effectively restored knowledge about famous people. The setup for connecting LoRA adapters is described in Appendix A.

### v2

@@ Abstract
The impressive capability of modern text-to-image models to generate realistic visuals has come with a serious drawback: they can be misused to create harmful, deceptive or unlawful content. This has accelerated the push for machine unlearning. This new field seeks to selectively remove specific knowledge from a model’s training data without causing a drop in its overall performance. However, it turns out that actually forgetting a given concept is an extremely difficult task. Models exposed to attacks using adversarial prompts show the ability to generate so-called unlearned concepts, which can be not only harmful but also illegal. In this paper, we present considerations regarding the ability of models to forget and recall knowledge, introducing the Memory Self-Regeneration task. Furthermore, we present MemoRa strategy, which we consider to be a regenerative approach supporting the effective recovery of previously lost knowledge. Moreover, we propose that robustness in knowledge retrieval is a crucial yet underexplored evaluation measure for developing more robust and effective unlearning techniques. Finally, we demonstrate that forgetting occurs in two distinct ways: short-term, where concepts can be quickly recalled, and long-term, where recovery is more challenging. Code is available at https://gmum.github.io/MemoRa/.
@@ 1 Introduction
Figure 2: An attempt to forget the concept of nudity by the Flux-based ESD model and the application of the MemoRa strategy. Results indicate that ESD has only temporarily forgotten this concept, making it possible to quickly recover it.
@@
Figure 3: Our method aims to recover unlearned information using only a few images that contain removed concepts. We first expand the training set using DDIM inversion and diversify it via spherical interpolation. Next, we fine-tune a LoRA adapter to restore the erased concept. Results reveal two types of forgetting: short-term, where knowledge is quickly recovered, and long-term, where recovery is harder. We hypothesize that short-term forgetting corresponds to superficial removal of knowledge, where concepts are not replaced but merely hidden, whereas long-term forgetting induces substantial changes in the distribution of data related to the forgotten concept, effectively altering the underlying representations.
@@
Figure 4: Visualizations of the MemoRa strategy applied to the FLUX.1 [dev] model for the "parachute" concept. Notably, MemoRa achieves a highly faithful restoration of the original visual characteristics, demonstrating precise recovery of the dormant knowledge.
@@
To investigate this phenomenon, we propose the MemoRa (Memory Regeneration with LoRA) strategy, which we consider a regenerative process, showing how knowledge recovery affects unlearning models. An overview of the strategy is presented in Fig. 3. Our method begins by applying DDIM inversion to reconstruct latent trajectories of images corresponding to the removed concept. To overcome the scarcity of available samples, we expand this dataset using spherical interpolation in latent space, which provides diverse yet consistent training examples. Next, instead of fine-tuning the full model, we update only a lightweight LoRA adapter, enabling efficient retraining with minimal computational cost. This design makes MemoRa practical even under resource constraints, while also serving as a diagnostic tool to probe the depth of forgetting. Some models fail to truly forget the targeted concept. Furthermore, we demonstrate that certain models exhibit a particular predisposition to rapidly recall previously forgotten concepts, indicating that this information is still stored in structures analogous to human long-term memory, we call this phenomenon short-term forgetting. This process is demonstrated on generated images using prompts containing concepts of nudity (Fig. 2), and parachute (Fig. 4). In contrast, models with lower regenerative capacity show greater memory loss and more effective unlearning.
Specific models are able to rapidly recover the erased knowledge, while others require extensive fine-tuning. We hypothesize that this difference arises from the way unlearning methods affect the underlying representation manifold. Our analysis indicates that the nature of forgetting is closely tied to the model’s trajectory relative to the data manifold, although the relationship is not straightforward. We find that Short-Term Memory (STM) forgetting reflects a form of functional suppression. Sometimes this appears as a brief decline in generation quality, pushing the model slightly off the manifold. In other cases, it acts as a superficial mask that keeps the model on the manifold while temporarily blocking access to the concept. In contrast, Long-Term Memory (LTM) forgetting tends to maintain strong alignment with the manifold by relocating the representation to a different semantic region. This shows that recoverability depends not only on optimization dynamics but also on the geometry of the shift. We provide a detailed analysis of these mechanisms and their geometric implications in Section 4.
@@
Figure 6: To enable recall without retraining, we augment limited samples using spherical interpolation. This respects the hyperspherical latent geometry, preserving concepts that linear interpolation often misses by deviating from the manifold.
@@ 2 Related Works
More recent work on unlearning within diffusion models focuses on parameter updates that suppress unwanted concepts. EraseDiff (ED) (Wu et al., 2024) achieves this through a bi-level optimization scheme, while ESD (Gandikota et al., 2023) modifies classifier-free guidance by incorporating negative prompts. FMN (Zhang et al., 2024a) introduces a targeted loss on attention layers to steer the forgetting process. SalUn (Fan et al., 2023) and SHS (Wu & Harandi, 2024) adapt model weights by exploiting saliency and sensitivity analyses to localize parameters relevant to the concept. SEMU (Sendera et al., 2025) leverages Singular Value Decomposition (SVD) to project representations into a lower-dimensional space that facilitates selective erasure. Selective Ablation (SA) (Heng & Soh, 2023) replaces the distribution of a forbidden concept with that of a surrogate, an idea extended in Concept Ablation (CA) (Kumari et al., 2023) using predefined anchors. In another direction, SPM (Lyu et al., 2024) employs structural interventions by inserting lightweight linear adapters that block the flow of targeted features. SAeUron (Cywiński & Deja, 2025) applies sparse autoencoders to isolate and suppress concept-specific representations, offering interpretable and robust unlearning with minimal performance degradation, even under adversarial prompting. On the other hand, AdvUnlearn (Zhang et al., 2024c) proposes using adversarial training to optimize the text encoder directly.
@@
There are various techniques for attacking DMs to bypass protections against malicious content generation. The basic methods involve manipulating words, which include replacing, deleting letters, inserting additional characters, and paraphrasing an unlearned concept (Eger & Benz, 2020; Li et al., 2018; Garg & Ramakrishnan, 2020). Other attack techniques include textual inversion (Gal et al., 2022), utilizing a frozen U-Net as a guide (white box) (Chin et al., 2023), or UnlearnDiffAtk (Zhang et al., 2024b), which generates adversarial prompts without needing an auxiliary model. The approaches discussed above enable the removal of targeted concepts from text-to-image models. Yet, regardless of the strategy employed, an important question remains: to what extent is the forgetting actually effective? Evidence suggests that current methods have notable limitations, as erased concepts can often be recovered. For instance, UnlearnDiffAtk (Zhang et al., 2024b) shows that both prompt tuning and backdooring techniques (Wang et al., 2024; Grebe et al., 2025) can restore removed concepts. Example images obtained from adversarial attacks are shown in Fig. 5. A key aspect of our pipeline, consistent with the methods discussed above, is the assumption of access to the model. It is worth noting that prompt tuning is computationally demanding, requiring separate attacks for each prompt, and fine-tuning alters concept representations in ways that make it difficult to pinpoint which concepts are affected. These studies highlight the inherent shortcomings of existing unlearning methods.
@@ 3 How to Relearn diffusion models using MemoRa strategy
Standard evaluation metrics primarily utilize a scheme for generating images from prompts that use words related to the unlearned phrase, paraphrase the concept, or use synonyms. This approach only analyzes the model’s performance in user-tested mode, but does not examine which representations remain and which are removed from the latent space. Fig. 7 displays the comprehensive inversion of the entire photo denoising process for two unlearned techniques. It investigates how different methods respond to encoding a prohibited image and evaluates the effectiveness of these methods in erasing the semantic features associated with the forbidden concept in the latent representation after inversion. If the images of this concept reproduce, we are dealing with short-term forgetting, while if the inversion process leads to deviated latencies relative to the original versions, we are dealing with long-term memory forgetting. It is noticeable that the ESD method is sensitive to repeated reconstruction of the deleted concept, whereas MACE is a more aggressive approach that removes this representation more efficiently.
@@
[…]
@@
Additionally, we use the extracted latents to generate more images by spherical interpolation, see Fig. 6. Spherical interpolation is a complement to Stage 1, as illustrated in Fig. 3, and determines intermediate values between two established latents.
@@
Figure 10: Reduced latent space representation of the I2P dataset for the nudity unlearning task using ED, ESD, and AdvUnlearn methods, alongside corresponding recall trajectories and FID comparisons. The ED method induces a substantial displacement from the original SD 1.4 manifold, leading to long-term forgetting and degraded images. Crucially, the FID analysis reveals a critical distinction between ESD and AdvUnlearn. While both methods exhibit similar levels of forgetting, they differ in FID values when compared to images generated using SD 1.4. This discrepancy indicates that the mechanism of forgetting varies fundamentally between them. It demonstrates that the geometric quality of the shift, rather than the absolute magnitude of the displacement, determines whether the forgetting corresponds to a reversible Short-Term state or a deeper Long-Term erasure.
4 Short- vs Long- term forgetting
We group the evaluated unlearning methods into two classes based on their recovery behavior: Short-Term (STM) and Long-Term (LTM) Memory forgetting. In the case of Nudity Concept Memory Recovery, STM is represented by SPM, ESD, and FMN. In contrast, LTM is represented by MACE, SalUn, AdvUnlearn, ED, and SH. As shown in Tab. 1, STM methods allow knowledge to be restored quickly, whereas LTM methods remain persistently resistant to recovery. Importantly, these differences arise from fundamentally distinct mechanisms rather than recovery speed alone.
@@ 4 Short- vs Long- term forgetting
Fig. 10 shows reduced latent space representation of the I2P dataset (Schramowski et al., 2023), in the task of unlearning nudity. We present the FID score between data generated by Stable Diffusion (before unlearning) and the outputs of unlearned models. We can see that the FID score is significantly higher for long-term than for short-term forgetting. We also visualized behaviors of unlearning in the reduced space using PCA technique.
To interpret these effects, we first consider their relationship to generation quality. When a method causes a substantial deterioration in FID (for example, ED), the model has likely drifted far from the natural image manifold. This is a clear form of LTM forgetting in which the model loses its ability to synthesize coherent images in general. In such cases, recovering the erased knowledge requires repairing the manifold itself, as illustrated in Fig. 10.
Method KRS ($\downarrow$)
Nudity Parachute Church Garbage Truck
FMN 41.53 91.67 75.00 48.15
UCE 18.92
SPM 57.82 64.86 75.00 20.83
ESD 60.18 80.85 72.09 48.98
MACE 17.06
SalUn 7.86 84.78 62.22 51.02
AdvUnlearn 18.36 4.08 18.00 8.00
ED 7.00 52.08 51.06 17.02
SH 0.00 0.00 2.00 0.00
Table 3: Knowledge Recovery Score (KRS) in percentages ($\%$) achieved by MemoRa across different unlearning methods for nudity and objects concepts. This metric allows for a hierarchical classification of methods relative to the MemoRa regeneration capabilities.
The comparison becomes more subtle for methods that maintain stable FID scores, such as ESD (STM) and AdvUnlearn (LTM). Although they differ significantly in recoverability, an analysis of the update vectors during regeneration shows that both require shifts of similar magnitude. This indicates that the optimization ,,distance” needed to restore the concept is comparable, yet the resistance to recovery diverges. We attribute this discrepancy to the underlying nature of the semantic change.
STM methods such as ESD appear to produce a prompt vs concept misalignment. In this regime, the visual features associated with the concept remain encoded in the model’s weights, but their linguistic triggers no longer reliably activate them. The model, therefore, ,,forgets” how to access the concept rather than the concept itself. Recovery is fast because self-regeneration does not need to reconstruct the visual representation. It only needs to remake the connection between the prompt and the intact latent features. This creates a superficial form of suppression in which knowledge remains hidden but still preserved.
In contrast, LTM methods such as AdvUnlearn induce Adversarial Overwriting. Here, the model is optimized against adversarial prompts forcing the parameters to reject the concept regardless of the prompt phrasing. As a result, self-regeneration must relearn the target features, facing a significantly more challenging optimization landscape.
In summary, although methods like ESD and AdvUnlearn may behave similarly under static metrics (see Tab. 1), they correspond to fundamentally different geometric states. STM reflects a reversible suppression in which knowledge is inaccessible yet preserved, whereas LTM reflects a stable semantic shift in which the original knowledge has been replaced.
5 Experiments
In this section, we present an evaluation of MemoRa for restoring knowledge across three main categories: nudity, objects, and styles, see Fig. 1. For this purpose, we utilized publicly available weights from unlearned models from Zhang et al. (2024c), which includes SOTA unlearning models. MemoRa is a strategy that stands out for its simplicity, making it easily applicable to all the aforementioned methods. Remarkably, MemoRa enables these models to recover lost knowledge in a much faster and straightforward manner. Additional experiments are presented in the Appendix.
@@ 5 Experiments
Evaluation Setups To assess the effectiveness and performance of MemoRa, we conduct a similar evaluation as in (Zhang et al., 2024c), which also includes testing the models’ robustness to attacks using adversarial prompts, employing the UnlearnDiffAtk technique. To evaluate MemoRa, we use a complex measure: the attack success rate (ASR) (Zhang et al., 2024b). To calculate effectiveness, a set of 50 prompts generated with GPT-4 was used, which contained target training words from the Imagenette dataset. These prompts were previously tested to ensure that the original Stable Diffusion 1.4 could generate correct images from them. The ASR metric can be divided into two components: the pre-attack success rate (pre-ASR) and the post-attack success rate (post-ASR). The pre-ASR metric reflects the model’s unlearned knowledge, as it is evaluated under normal conditions without any attacks. In contrast, the post-ASR metric assesses the effectiveness of attacks on the unlearned model using adversarial prompts. These metrics are not correlated, as high unlearned data will not always lead to low attack effectiveness. Therefore, for the purposes of our experiment, we consider these two metrics separately. The "No Attack" scenario will represent the pre-ASR, and "UnlearnDiffAtk" will encompass both pre-ASR and post-ASR to allow for a more thorough assessment of the relearning of the models mentioned. The higher the "No Attack" measure, the better the model generates images in default mode. Conversely, a higher "UnlearnDiffAtk" measure indicates lower resistance to attacks. We used ImageNet-pretrained ResNet-50 to classify images for attack evaluation.
Method Train Set Eval Set Armpits Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M)
ESD Women Men 137(38) 95(30) 1(3) 11(12) 39(10) 17(4) 117(22) 22(1)
Men Women 127(35) 91(19) 9(7) 2(12) 158(39) 49(5) 0(3) 1(1)
UCE Women Men 99(38) 88(24) 1(1) 7(0) 69(18) 1(3) 75(30) 10(3)
Men Women 132(50) 92(29) 1(1) 5(0) 167(95) 63(7) 2(4) 0(0)
SD v1.4 Men 109 80 4 6 37 6 90 24
Women 108 76 13 4 158 44 0 0
Table 4: Cross-Gender Memorization. Experiments for MemoRa, which uses single-gender images ("a photo of the nude man" or "a photo of the nude woman") to remind the concept of nudity for the opposite gender. Detection of unsafe body parts using the NudeNet classifier for 100 images, focusing on categories specifically related to gender: women (F) and men (M). Ultimately, MemoRa reinstates nudity overall, aligning it more closely with the original distribution across all classes. The values in brackets are the results for the unlearned model.
Config Type Variant Metrics
No-Attack $(\uparrow)$ FID $(\downarrow)$ CLIP $(\uparrow)$
1 Pairs 1 83.05% 31.89 29.86
3 68.31% 22.57 30.58
6 66.95% 22.07 30.63
2 $t$ 45 58.47% 23.23 30.51
40 62.71% 23.62 30.56
35 68.31% 22.57 30.58
SD v1.4 100% 17.02 31.08
Table 5: Ablation study of the key components the nudity dataset creation. Pairs: the number of image pairs used for inversion and interpolation; $t$: the time step at which the inversion process stops.
@@
To quantify how effectively MemoRa restores forgotten knowledge, we introduce the Knowledge Recovery Score (KRS) calculated as follows:
$\text{KRS}=\frac{\text{Perf}(\text{MemoRa})-\text{Perf}(\text{Unlearned})}{\text{Perf}(\text{Base})-\text{Perf}(\text{Unlearned})}$
where Perf() is the accuracy of the classifier (e.g. ResNet, NudeNet Detector) that checks whether the generated image contains a specific concept.
@@
Nudity Relearning Tab.1 presents the results using the MemoRa strategy. Each method experienced an increase in knowledge, but at varying levels. MemoRa suggests that SPM and ESD are methods in which unlearning is shallow and a large return to pre-unlearning knowledge is possible. MemoRa allowed for the unlearned knowledge to be recovered, as symbolized by the significantly increasing No Attack measure. Tab. 3 shows the KRS scores calculated based on the No Attack metric for nudity restoration. Fig. 8 compares all unlearning methods that used the MemoRa strategy. The SH, SalUn, and ED methods remove nudity features the most in the generated images even after applying the memory regeneration strategy. The disadvantage of these approaches is the deterioration of the remaining knowledge, where MemoRa corrects it closer to the original state. Sample visualizations of this situation are presented in Fig. 9. More visual and quantitative results for MemoRa are provided in the Appendix B.1.
@@
Objects Relearning Three classes from the Imagenette (Shleifer & Prokop, 2019) dataset were used: parachute, church and garbage truck, see Fig. 5. For objects, it is evident that some models forget more while others forget less, as illustrated by the KRS metric in Table 3. MemoRa increases the number of successful attacks, leading to shorter inference times using adversarial prompts, see Tab. 2. Introducing attacking involves time-consuming inference, whereas with MemoRa + UnlearnDiffAtk strategy, inference time is much shorter (for example from 10 minutes to 4 minutes). MemoRa requires only pre-training (approximately 15 minutes), and its image generation time matches that of the base model. More visual and quantitative results are provided in the Appendix B.2.
MemoRa on FLUX. 1 The MemoRa strategy was tested on the FLUX.1 [dev] with inference step as 28 for the nudity and parachute concepts. Only ESD and UCE were adopted as unlearning models because they are compatible with the FLUX pipeline. The RF-Inversion (Rout et al., 2025) was employed for the inversion process.
Method Results of NudeNet Detection
Armpits Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M) No Attack $(\uparrow)$
ESD 15 17 0 4 16 0 4 0 53.85%
ESD + MemoRa 43 19 0 6 22 0 9 0 94.24%
FLUX. 1 [dev] 42 25 2 8 26 1 6 0 100.0%
Table 6: Evaluation of Nudity Concept Memory Recovery for FLUX.1 [dev]-ESD model. Amount of explicit content found using the NudeNet detector on the I2P dataset. The last column shows the average percentage of detected examples compared to FLUX.1[dev] model.
Model Variant Metrics
No-Attack $(\uparrow)$ FID $(\downarrow)$ CLIP $(\uparrow)$
ESD + MemoRa linear 66.95% 24.45 30.39
ESD + MemoRa spherical 68.31% 22.57 30.58
UCE + MemoRa linear 36.44% 18.63 30.93
UCE + MemoRa spherical 36.62% 18.57 30.98
SD v1.4 100% 17.02 31.08
Table 7: Numerical comparison of linear vs. spherical latent interpolation for the task of restoring the Nudity concept. Linear interpolation produces worse quality results than spherical interpolation.
Cross-Gender Evaluation An experiment was conducted that confirms that MemoRa enables the model to recall knowledge instead of simply fine-tuning on a specific training set. A detailed comparison of ESD and UCE is shown in Fig. 4, which reveals that, although we used only images of naked men to restore the nudity concept, the features of naked women were also significantly restored.
Spherical vs. Linear Interpolation Effects To study the impact of the type of interpolation in preparing the dataset, a comparison was made between linear (LERP) and spherical (SLERP) interpolation for restoring the nudity concept. Fig. 6 illustrates the complete 10-step intermediate process of interpolation for latents in the ESD model. LERP tends to produce unnatural transitions, resulting in distorted or blurry images. In contrast, SLERP creates smooth transitions while preserving realistic and semantically consistent images. Tab.7 presents the numerical results of this comparison, showing that SLERP does not lead to a sudden increase in the FID measure.
Effect of Pairs and Timestep $t$ A comparison of results for various numbers of training images and different timesteps $t$ was presented in Tab.5. For 3 image pairs, the model successfully recovers a significant portion of the knowledge, demonstrating stable CLIP and FID results. Notably, increasing the number of pairs to 12 does not lead to a proportional improvement. This shows that the model retains information well, so few examples are enough to access the forgotten knowledge effectively. The optimal step size for stopping the inversion was tt=35, as it balances latent noise with the information required for the unlearned model to effectively reproduce the target concept.